A pipe 0.0303 m in diameter and 0.43 m long can tr
Answer #1


H = KA(T2 - T1)/x

H = 427*(3.14*0.0303^2/4)(6)/0.43

H = 4.5 W

For solid silver Q = 4.5 J

Qgiven = 3543 J

Qgiven/Q = 3543/4.5 = 787.3